### v1

# Study on physical properties and characteristics of an anisotropic compact star model using Karmarkar Condition in $\mathcal{F}(\mathcal{Q})$ gravity

Abstract
The main aim of this study is to examine the behaviour of physical parameters of an anisotropic compact star model demonstrating spherical symmetry in $\mathcal{F}(\mathcal{Q})$ modified gravity. To evaluate the behaviour and the stability of an anisotropic compact star model, we utilise the measured mass and radius of an anisotropic compact star model. This study obtained an anisotropic compact star model by solving Einstein’s field equations. The field equations have been simplified by an appropriate selection of the metric element $e^{\sigma}$ and the Karmarkar condition. By solving field equation to develop a differential equation that establishes a relationship between two essential components of spacetime, namely $e^{\sigma}$ and $e^{\Upsilon}$. A physical analysis of this model reveals that the resulting stellar structure for anisotropic matter distribution is a physically plausible representation of a compact star with an energy density of order $10^{14}g/cm^{3}$. Using the Tolman-Oppenheimer-Volkoff equation, causality condition and Harrison-Zeldovich-Novikov Condition, we investigate the hydrostatic equilibrium and stability of the compact star Cen X-3. We further determined the mass-radius relation of this compact star for different values of $\delta_{1}$.
Keywords: compact star, anisotropic fluid sphere, modified gravity theory, Karmarkar condition
I Introduction
Various universal events  SupernovaSearchTeam:1998fmf ; Planck:2015mvg ; Planck:2015bpv have demonstrated that improvements to general relativity (GR) are essential at a geometric scale. GR, grounded on Riemannian geometry, illustrates the Levi-Civita connection and defines the Ricci curvature $R$ as the essential characteristic of spacetime. This approach omits torsion and non-metricity from the geometry. An alternate approach to characterise GR and Riemannian geometry is teleparallel gravity  aldrovandi2012teleparallel . This theory suggests that the gravitational force originates from torsion $\mathcal{T}$ rather than curvature $R$ in Riemannian geometry. This idea was formally proposed by Einstein, who had earlier utilised teleparallel geometry in his efforts to formulate a unified field theory  einstein2005riemann . A different technique, referred to as non-metricity or symmetric teleparallel gravity, was developed by Nester  nester1998symmetric . This concept suggests that non-metricity $\mathcal{Q}$, which is not impacted by curvature and torsion, acts as the mediator of gravitational interactions. In contrast to General Relativity, where gravity is conveyed through the affine connection instead of the physical manifold, non-metricity is essential in this alternative approach. The Einstein pseudo-tensor, which transforms into a genuine tensor in the geometric framework, is associated with the non-metricity $\mathcal{Q}$. Jimenez et al. BeltranJimenez:2017tkd expanded symmetric teleparallel gravity to align with general relativity, obtaining $F(\mathcal{Q})$ gravity. Given the growing focus on extended gravity theories in modern cosmology, researchers are exploring alternate geometries to achieve a deeper understanding of late cosmic acceleration  Ferraro:2008ey ; Geng:2011aj ; Cai:2015emx ; Jarv:2015odu . Furthermore, it is important to acknowledge that the metric theories may be expanded beyond Riemannian geometry, as shown in  Conroy:2017yln . Harko et al. Harko:2018gxr presented the modified $F(\mathcal{Q})$ gravity by utilizing the metric-affine formalism and an expansion of symmetric teleparallel gravity. Recent research on advanced theories of gravity emphasises the critical examination of the astrophysical and cosmological characteristics of $F(\mathcal{Q})$ gravity. In this way, Hohmann et al. Hohmann:2018wxu investigated the propagation velocity and possible polarizations of gravitational waves within the framework of Minkowski spacetime in $F(\mathcal{Q})$ gravity. Soudi et al. Soudi:2018dhv defined the strong-field characteristics of this gravitational theory by the examination of gravitational wave polarizations. A multitude of studies has investigated $F(\mathcal{Q})$ gravity across diverse contexts, including the analysis of observational data constraints for various parametrizations of $F(\mathcal{Q})$  Lazkoz:2019sjl ; Ayuso:2020dcu , energy conditions  Mandal:2020lyq , cosmography  Mandal:2020buf , bouncing scenarios  Mandal:2021wer ; bajardi2020bouncing , black holes DAmbrosio:2021zpm , and the development of the growth index in matter perturbations, among other issues Nguyen:2021cnb . Furthermore, comprehensive analyses of the $F(\mathcal{Q})$ theory may be found in contemporary literature, including Esposito:2021ect ; Atayde:2021pgb ; Kimura:2020mpk ; Dimakis:2021gby . It is crucial to highlight that $F(\mathcal{Q})$ gravity precisely accounts for the universe’s fast expansion with statistical accuracy comparable to other established modified gravity theories. Therefore, altered gravity theories at an astrophysical scale must be taken in order to deal with this so-called degeneracy. Promising approaches for this purpose include direct monitoring of the galactic core  Mann:2018xkm and the increasing number of gravitational wave findings  LIGOScientific:2016aoc . For a comprehensive examination of this subject, the subsequent sources  errehymy2022anisotropic ; Maurya:2022cyv ; Tangphati:2021tcy include modern contributions across several settings. Stellar evolution describes the ongoing alterations in a star’s interior composition from its starting point to its final extinction. During this transforming process, a star releases a significant quantity of energy as photons or neutrinos. The fascinating properties of these celestial bodies have inspired many scholars to investigate their significance in astronomy and astrophysics. Gravitational collapse is a crucial topic in relativistic astrophysics, intimately linked to the formation of self-gravitating objects. A stable system initially sustains hydrostatic equilibrium until its gravitational forces continuously decrease its size, leading to gravitational collapse. The collapse yields many configurations, including white dwarfs, neutron stars, black holes and gravastars, called compact stars contingent upon the star’s starting mass. The exact solution of an innovative gravastar model derived from the Mazur:2004fk approach within the context of general relativity, particularly through incorporating the cloud of strings and quintessence in paper Javed:2023nls . The composition and features of compact stars have fascinated scientists, leading to inquiries into their many developmental phases and intrinsic properties within the field of gravitational physics. Multiple endeavours have been undertaken to examine the properties of stellar models influenced by diverse physical forces. In this context, the concept of anisotropy is essential for examining the structural processes of compact structures  Herrera:1997plx . Bower and Liang  bondi1992anisotropic examined regionally anisotropic pressure distributions inside spherically symmetric matter, demonstrating that pressure anisotropy significantly influences the parameters defining the hydrostatic equilibrium of stellar systems. Maurya and his associates Maurya:2017uyk . examined compact spherically symmetric systems characterized by anisotropic pressure profiles, providing substantial insights into the stability analysis of small objects. Karamakar and colleagues  paul2015relativistic investigated the influence of anisotropy in cool compact stars. A large number of scholars have exhibited a keen interest in examining the intrinsic and extrinsic characteristics of celestial bodies using gravitational elements. The Karmarkar condition is a crucial framework that establishes a relationship between temporal and radial metric elements, hence improving the examination of compact objects. The Karmakar condition deals with the embedding of a four-dimensional Riemannian manifold into a higher-dimensional manifold. This approach is utilized to investigate the solutions of static spherical spacetime and to analyze the physically viable characteristics of anisotropic stars  Maurya:2016ecj ; Singh:2016nfw . Bhar et al.  Bhar:2017jbl used this technique to investigate the stability requirements of anisotropic compact stars.
Uneven primary stresses or the so-called anisotropic fluid are expected to appear when the densities of compact objects are typically higher than the density of nuclear matter. This often indicates that these compact objects have two distinct types of pressures: the tangential pressure $(p_{t})$ and the radial pressure $(p_{r})$ Herrera:1997plx . The radial pressure component $(p_{r})$ is not equal to the components in the transverse direction $(p_{t})$, resulting in the anisotropic condition. J.H. Jeans initially predicted this effect for self-gravitating objects in the Newtonian regime in 1922 Hueckstaedt:2006bt . Shortly after, Lemaitreluminet2015lemaitre had also taken into account the local anisotropy effect in the context of GR and shown that the higher limits placed on the maximum value of the surface gravitational potential might be relaxed. Ruderman ruderman1972pulsars provided an intriguing overview of more realistic stellar models and demonstrated that an anisotropic star is likely to have a matter density ($\rho>10^{15}gm/cm^{3}$), where the nuclear interaction takes on a relativistic character. Additionally, it is evident from Herrera’s recent explanation Herrera:2020gdg that even if the starting configuration had isotropic pressure, it would vanish and the system would become anisotropic owing to energy dissipation in stellar growth. However, a system has no cause to lose its acquired anisotropy in the last stage of dynamic development if the matter distribution is already anisotropic. One of the main characteristics of the growth of massive stars is this dissipation, which is caused by the emission of massless particles like photons or neutrons.
Makalo et al.  Makalo:2023anc examined exact solutions for charged stars employing the Karmarkar condition and found that the relationship between metric components and charge leads to the formulation of a viable anisotropic model. Sharif and Gul  gul2024physical investigated physically feasible and stable anisotropic stellar models in the context of the Karmarkar framework. The physical characteristics of stellar objects have been analyzed in several modified gravitational theories utilising the Karmarkar condition  Abbas:2019hyp ; Sharif:2022cjv ; hassan2023decoupled . Spherically symmetric solutions are examined extensively within the context of $F({R})$ gravity Multamaki:2006zb ; delaCruz-Dombriz:2009pzc ; Kobayashi:2008tq ; Upadhye:2009kt ; Manzoor:2022fjd . The same symmetry is also examined inside the $F(\mathcal{T})$ framework  Aftergood:2014wla ; DeBenedictis:2016aze ; lin2017spherical ; Krssak:2015oua . The dynamic characteristics of compact stars have been examined in several alternative theories of gravity, such as $F(\mathcal{G})$  Abbas:2014vka ; malik2022bardeen , $F(\mathcal{Q})$  Mandal:2021qhx ; Errehymy:2022gws , and $F(\mathcal{G,T})$  Shamir:2017rjz ; AwaisSadiq:2022pwn , among others. Numerous other investigations on compact stars exist within various modified theories and scenarios Bhar:2017jbl ; Malik:2024eui ; Malik:2024boe ; Malik:2024vjm ; Malik:2024nfl ; Malik:2024tcg ; Ashraf:2024cww ; Bhar:2020tah ; Bhar:2021fzc ; Bhar:2017ynp .
Narawade et al. in Narawade:2024pxb reconstructs a cosmological model utilizing covariant $F(\mathcal{Q})$ gravity within Connection-III and FLRW spacetime. The dynamic behaviour of the reconstructed model is thoroughly analyzed using the H(Z) Hubble parameter and various observational datasets. This research offers compelling insights into cosmic acceleration, suggesting that $F(\mathcal{Q})$ gravity can effectively displace the $\Lambda CDM$ model and provides a convincing alternative explanation for the current accelerating expansion of the Universe without relying on the cosmological constant. Using a dynamical system method, Lohakare et al. Lohakare:2024oeu investigated the cosmic stability of $F(\mathcal{Q,B})$ gravity, where $F(\mathcal{Q})$ stands for the nonmetricity scalar and $F(\mathcal{B})$ for the boundary term. According to this study, the $F(\mathcal{Q,B})$ model is a good substitute for the conventional $\Lambda CDM$ model as it provides a strong framework for understanding the dynamics of cosmic expansion and accurately depicts the observed acceleration of the Universe. In  Alwan:2024lng Alwan et al. investigate neutron stars with covariant $F(\mathcal{Q})$ gravity, assuming that the stars are spherically symmetric and static with ideal fluid matter. Numerical techniques employing quadratic, exponential, and logarithmic $F(\mathcal{Q})$ models are used to determine neutron star characteristics including mass, radius, and compactness. A cosmological model within the context of symmetric teleparallel gravity is presented by Narawade et al.in  Narawade:2023rip . They discover a violation of the strong energy requirement, which may offer important insights into the nature of dark energy. In  Narawade:2023tnn , the accelerating cosmological model’s dynamical aspect was examined. Additionally, the model’s evolutionary behaviour of density parameters for the matter-dominated, radiation-dominated, and dark energy phases is demonstrated within the framework of the $F(\mathcal{Q})$ gravity, a modified symmetric teleparallel gravity. In Narawade:2022cgb , a cosmological model of the Universe is presented within the framework of $F(\mathcal{Q})$ gravity, and it is shown that $F(\mathcal{Q})$ gravity offers an alternative to dark energy for explaining the current cosmic acceleration.
In the present study, the main focus was on $F(\mathcal{Q})$ gravity. The modified gravity provides a straightforward composition of the classical GR, which contributes to improving inertial gravitational interaction, without taking into accounting for the affine space-time structure. The special feature of $F(\mathcal{Q})$ gravity is its capacity to interpret the Universe’s late-time acceleration without demanding alterations to scalar fields. Akarsu:2010zm ; BeltranJimenez:2019tme ; Lazkoz:2019sjl . The viability of $F(\mathcal{Q})$ gravity model has been proved in a wide range of domains, including wormhole geometry Harko:2018gxr ; Xu:2019sbp ; DAmbrosio:2020nev ; Maurya:2022vsn ; Maurya:2023kjx ; Banerjee:2021mqk ; Kiroriwal:2023nul and dynamical analysis Lu:2019hra , which covers the cosmological aspects Mandal:2020buf , and has been presented in relation to $F(\mathcal{Q})$ gravity. The mathematical representation of $F(\mathcal{Q})=\mathcal{Q}+m\mathcal{Q}^{n}$ has been used to investigate holographic dark energy Shekh:2021ule which reveals the behaviour of the deceleration parameter changes from negative to positive, indicating that the universe is transitioning from an early phase of deceleration to the current phase of acceleration. In the framework of extended non-metric theories of gravity, Capozziello et al.  Capozziello:2022tvv explored the possibility of cosmological slow-roll inflation. Specifically, they found that standard conformal structure is recovered in a high-energy regime while it breaks in the weak field limit, as demonstrated by generic functions $F(\mathcal{Q})$, particularly $F(\mathcal{Q})=\alpha Q+\beta Q^{m}$. Regardless of the extensively accepted, the teleparallel version of GR shares a formulation with both STGR and GR  Heisenberg:2022mbo . The primary contrast lies in that gravity is allocated to the non-metricity $(\mathcal{Q})$ , whereas curvature and torsion are regarded as zero in STGR. Compatible with the non-metricity and zero value of the torsion restriction, a coincident gauge is always applicable BeltranJimenez:2017tkd . This method is unique for the reason that it eliminates the affine connections. Nevertheless, the development of metrics will differ in different coordinate systems if we continue with the coincident gauge Lu:2019hra and treat the metric as the only fundamental variable in modified STGR theories, such as the $F(\mathcal{Q})$ theory. While research on extended gravity theories is still in its early stages, the symmetric teleparallel, or $F(\mathcal{Q})$, theory of gravity has been receiving a lot of attention. Due to these novel features of $F(\mathcal{Q})$ gravity, several studies have been conducted so far to examine various aspects of the cosmos, including cosmology, energy requirements, covariant formulation, spherical symmetric structure, and signatures that highlight the significance of gravity. The idea behind working in the $F(\mathcal{Q})$ theory of gravity is that on comparing to the $F(\mathcal{R})$ theory, which involves differential equations of the fourth order, the field equations of $F(\mathcal{Q})$ gravity are of the second order Lin:2021uqa , $F(\mathcal{Q})$ gravity extends the theoretical foundation for modelling compact stars by giving new approaches to examining their structure, stability, and observational signature. It offers a promising path for addressing issues concerning severe astrophysical conditions and challenging the limits of our understanding of gravity which is a significant advantage.
This study presents a novel anisotropic compact star model in $F(\mathcal{Q})$ gravity derived from the resolution of Karmarkar condition in static spherical spacetime. There exist two categories of solutions to the Karmarkar condition one is the interior Schwarzschild solution (1916) possesses an inaccurate causality condition and the other is cosmic. In our research, we use a particular kind of metric potential characterized by a hyperbolic function. This is feasible within the physical constraints of the astrophysical model. Developing solutions can be utilized to characterize a physically plausible distribution of astrophysical matter. We have conducted a comprehensive analysis of all physical characteristics and provide sample figures to substantiate our results. The structure of the article is as follows:
Section I: Introduction
Section II: In this section, we have discussed the field equations and the Karmakar condition.
Section III: In this section, we explain the physical properties and characteristics of our proposed model.
Section IV. This section contains the conclusion part.
II Field equations of $\mathcal{F}(\mathcal{Q})$ gravity and Karmarkar condition
II.1 Field equations of $\mathcal{F}(\mathcal{Q})$ gravity
The Levi-Civita affine connection is crucial in explaining and analysing general relativity. When discussed inside a space-time manifold, it aligns with the metric. Various manifolds and affine connections may be utilized, along with diverse theories of gravity. Harada:2020ikm ; BeltranJimenez:2019tme .The Levi-Civita connection identifies nonmetricity $\mathcal{Q}$ and torsion $\mathcal{T}$ as important geometrical concepts alongside curvature. Alleviating these limitations permits the exploration of non-Riemannian geometric theories characterized by non-vanishing curvature, torsion, and nonmetricity. The $\mathcal{F}(\mathcal{Q})$ gravity model was formulated under these circumstances, yielding a gravitational Lagrangian that incorporates an arbitrary nonmetric function ($\mathcal{Q}$). The extension of $\mathcal{F}(\mathcal{Q})$ gravity is extremely significant in the expansion of the Universe. Jimenez et al BeltranJimenez:2019tme was the first to coin the phrase ”symmetric teleparallel gravity,” commonly known as $\mathcal{F}(\mathcal{Q})$ gravity. Based upon the research of  Zhao:2021zab within the framework of $\mathcal{F}(\mathcal{Q})$ gravity, a generic metric spacetime is analyzed, in which the metric tensor $g_{\gamma\Upsilon}$ and the connection coefficients $\Gamma^{\sigma}_{\gamma\Upsilon}$ are treated as distinct entities.The subsequent equation delineates the Nonmetricity of the aforementioned connection:
$\mathcal{Q}_{\psi\gamma\Upsilon}=\nabla_{\psi}g_{\gamma\Upsilon}=\partial_{\psi}g_{\gamma\Upsilon}-\Gamma^{\sigma}_{\psi\gamma}g_{\sigma\Upsilon}-\Gamma^{\sigma}_{\psi\Upsilon}g_{\gamma\sigma}.$ (1)
These three components, in their generic form, are obtained by disassembling this affine relationship into independent components:
$\Gamma^{\sigma}_{\gamma\Upsilon}=\{^{\sigma}{}_{\gamma\Upsilon}\}+\zeta^{\sigma}_{\gamma\Upsilon}+L^{\sigma}_{\gamma\Upsilon},$ (2)
where $\{^{\sigma}{}_{\gamma\Upsilon}\}$ denotes the levi-Civita connection, $\zeta^{\sigma}_{\gamma\Upsilon}$ denote the Contortion tensor and $L^{\sigma}_{\gamma\Upsilon}$ denotes the disformation. $\{^{\sigma}{}_{\gamma\Upsilon}\}$ is solved by using the metric potential $g_{\gamma\Upsilon}$ :
$\{^{\sigma}{}_{\gamma\Upsilon}\}\equiv\frac{1}{2}g^{\sigma\Psi}(\partial_{\gamma}g_{\Psi\Upsilon}+\partial_{\Upsilon}g_{\Psi\Upsilon}-\partial_{\Psi}g_{\gamma\Upsilon}),$ (3)
$\zeta^{\sigma}_{\gamma\Upsilon}$ is written as:
$\zeta^{\sigma}_{\gamma\Upsilon}\equiv\frac{1}{2}\mathcal{T}^{\sigma}_{\hskip 2.84544pt}{\gamma\Upsilon}+\mathcal{T}_{(\gamma\hskip 8.5359pt\Upsilon)}^{\hskip 5.69046pt\sigma}$ (4)
and $L^{\sigma}_{\gamma\Upsilon}$ is written as:
[…]
The nonmetricity conjugate is defined as:
$H^{\psi}_{\gamma\Upsilon}=\frac{-1}{4}\mathcal{Q}^{\psi}_{\gamma\Upsilon}+\frac{1}{2}\mathcal{Q}^{\psi}_{(\gamma\Upsilon)}+\frac{1}{4}(\mathcal{Q}^{\psi}-\tilde{\mathcal{Q}}^{\psi})g_{\gamma\Upsilon}-\frac{1}{4}\delta^{\psi}_{(\gamma\Upsilon)}.$ (6)
The independent traces of the preceding equations are as follows:
[…]
now, the nonmetricity scalar is defined as:
$\mathcal{Q}=-\hskip 2.84544pt\mathcal{Q}_{\psi\Upsilon\gamma}F^{\psi\gamma\Upsilon}.$ (8)
The $\mathcal{F}(\mathcal{Q})$ gravity is defined by the following action using Lagrange multiplets.:
$S=\int\sqrt{-g}d^{4}x\Bigg{[}\frac{1}{2}\mathcal{F}(\mathcal{Q})+\sigma^{\Psi\gamma\Upsilon}_{\psi}\mathcal{R}^{\psi}_{\Psi\gamma\Upsilon}+\sigma^{\gamma\Upsilon}_{\psi}\mathcal{T}^{\psi}_{\gamma\Upsilon}+\mathcal{L}_{n}\Bigg{]}.$ (9)
In Eq. (9), $g$ represents the determinant metric, $\mathcal{F}(\mathcal{Q})$ denotes the nonmetricity of $\mathcal{Q}$, $\mathcal{L}_{n}$ represents the Lagrange density, and $\sigma_{\psi}^{\Psi\gamma\Upsilon}$ represents the Lagrange multipliers. The action given in Eq. (9) has been used to derive the field equations concerning the metric given below
$\mathcal{T}_{\gamma\Upsilon}=\frac{2}{\sqrt{-g}}\Delta_{\psi}(\sqrt{-g}\mathcal{F}_{\mathcal{Q}}H^{\psi}_{\gamma\Upsilon})$ (10)
$+\frac{1}{2}g_{\gamma\Upsilon}\mathcal{F}+\mathcal{F}_{\mathcal{Q}}(H_{\gamma\psi\Psi}\mathcal{Q}_{\Upsilon}^{\psi\Psi}-2\mathcal{Q}_{\psi\Psi\gamma}H^{\psi\Psi}_{\Upsilon}).$
In equation Eq. (10), the subscript $\mathcal{Q}$ denote the derivative of the function $\mathcal{F}(\mathcal{Q})$ with respect to $\mathcal{Q}$, the energy momentum tensor is defined as :
$\mathcal{T}_{\gamma\Upsilon}\equiv\frac{2}{\sqrt{g}}\frac{\delta(\sqrt{-g})\mathcal{L}_{n}}{\delta g^{\gamma\Upsilon}}.$ (11)
We get the following equations when we change equation Eq. (9) concerning affine connection:
$\nabla_{\rho}\sigma_{\psi}^{\Upsilon\gamma\rho}+\sigma_{\psi}^{\gamma\Upsilon}=\sqrt{-g\mathcal{F}_{\mathcal{Q}}}H_{\gamma\Upsilon}^{\psi}+J_{\psi}^{\gamma\Upsilon}.$ (12)
The hypermomentum tensor density is represented by the following equation:
$J_{\psi}^{\gamma\Upsilon}=\frac{-1}{2}\frac{\delta\mathcal{L}_{n}}{\delta T_{\gamma\Upsilon}^{\psi}}.$ (13)
To simplify equation Eq. (12) to the following equation, we use the anti-symmetry property of s and t in the lagrangian multiplier coefficients:
$\nabla_{\gamma}\nabla_{\Upsilon}(\sqrt{-g}\mathcal{F}_{\mathcal{Q}}H^{\gamma\Upsilon}_{\hskip 8.5359pt\psi}+J_{\psi}^{\hskip 8.5359pt\gamma\Upsilon})=0.$ (14)
[…]
$\nabla_{\gamma}\nabla_{\Upsilon}(\sqrt{-g}\mathcal{F}_{\mathcal{Q}}H^{\gamma\Upsilon}_{\hskip 8.5359pt\psi})=0.$ (15)
The affine connection without torsion or curvature has the following form:
$\Gamma^{\psi}_{\gamma\Upsilon}=\Bigg{(}\frac{\partial{x}^{a}}{\partial{\eta}^{\sigma}}\Bigg{)}\partial_{\gamma}\partial{\Upsilon}\eta^{\sigma}.$ (16)
If we consider a specific coordinate selection that is well-known as a coincident gauge, for which $\Gamma^{\psi}_{\gamma\Upsilon}=0$. After that, the nonmetricity becomes:
$\mathcal{Q}_{\psi\gamma\Upsilon}=\partial_{\psi}g_{\gamma\Upsilon}.$ (17)
The metric for a general static and spherically symmetric space is expressed as :
$ds^{2}=-e^{\Upsilon(r)}dt^{2}+e^{\sigma(r)}dr^{2}+r^{2}dw^{2}.$ (18)
In the above equation $dw^{2}\equiv d\theta^{2}+sin^{2}{\theta}d{\phi}^{2}$. we get Eq. (19) by putting Eq. (18) in Eq. (8), then $\mathcal{Q}$ is expressed in terms of r,
$\mathcal{Q}(r)=-\frac{2e^{-\sigma}}{r}\Bigg{(}\Upsilon^{\prime}+\frac{1}{r}\Bigg{)},$ (19)
where prime($\prime$) signifies the derivative with respect to r. An anisotropic fluid with spherically symmetric geometry has the following energy-momentum tensor:
$\mathcal{T}_{\gamma\Upsilon}=(\rho+p_{t})u_{\gamma}u_{\Upsilon}+p_{t}g_{\gamma\Upsilon}+(p_{r}-p_{t})v_{\gamma}v_{\Upsilon},$ (20)
where, $u_{\gamma}$ denotes four-velocity, and $v_{\gamma}$ denotes the unitary space-like vector in the radial direction that fulfils $u^{\gamma}u_{\gamma}=-1,v^{\gamma}v_{\gamma}=1,u^{\gamma}u_{\gamma}=-1,u^{\gamma}v_{\Upsilon}=0$. $\rho(r)$ represents energy density, $p_{r}(r)$ represents the radial pressure in the direction of $v_{\gamma}$, while the orthogonal represents the tangential pressure $(v_{\gamma})$ is denoted by $p_{t}(r)$.The following are the independent components of the equations of motion Eq. (10) for anisotropic fluid Eq. (12) are written as follows:
$\rho=-\frac{\mathcal{F}}{2}+\mathcal{F}_{\mathcal{Q}}\Bigg{[}\mathcal{Q}+\frac{1}{r^{2}}+\frac{e^{-\sigma}}{r}(\Upsilon^{\prime}+\sigma^{\prime})\Bigg{]},$ (21)
$p_{r}=\frac{\mathcal{F}}{2}-\mathcal{F}_{\mathcal{Q}}\Bigg{[}\mathcal{Q}+\frac{1}{r^{2}}\Bigg{]},$ (22)
$p_{t}=\frac{\mathcal{F}}{2}-\mathcal{F}_{\mathcal{Q}}\frac{\mathcal{Q}}{2}-\mathcal{F}_{\mathcal{Q}}e^{-\sigma}\frac{\mathcal{Q}}{2}$
$-\mathcal{F}_{\mathcal{Q}}e^{-\sigma}\Bigg{[}\frac{\Upsilon^{\prime\prime}}{2}+(\frac{\Upsilon^{\prime}}{4}+\frac{1}{2r})(\Upsilon^{\prime}-\sigma^{\prime})\Bigg{]},$ (23)
$0=\frac{\cot\theta}{2}\mathcal{Q}^{\prime}\mathcal{F}_{\mathcal{Q}\mathcal{Q}}.$ (24)
Assuming a zero affine connection in the coordinate system, $\mathcal{F}(\mathcal{Q})$ gravity Eq. (9) can be expressed as follows:
$\frac{\cot\theta}{2}\mathcal{Q}^{\prime}\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0.$ (25)
Combining equation (10) and motion equations yields $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$. This leads us to the conclusion that the function $\mathcal{F}(\mathcal{Q})$ must be linear. Choosing nonlinear values for the function $\mathcal{F}(\mathcal{Q})$ might result in conflicting equations of motion and solutions, especially for $\mathcal{F}(\mathcal{Q}=\mathcal{Q}^{2})$. Applying a nonlinear function of $\mathcal{F}(\mathcal{Q})$ in gravity results in different values for the equation of motion. To analyze and resolve the nonlinear function of $\mathcal{F}(\mathcal{Q})$ gravity, we need a more comprehensive spherically symmetric metric for coincident gauges. Please visit  Zhao:2021zab for a more in-depth study of this topic. In this investigation, we used a linear function with $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$ and found that the spherically symmetric coordinate system (18) corresponds to the affine connection $\Gamma^{\gamma}_{\gamma\Upsilon}=0$. To formulate a realistic compact stellar model, we need to discover a useful function of $\mathcal{F}(\mathcal{Q})$. According to  Wang:2021zaz , matching the compact star solution with the Schwarzchild (Anti-) de sitter solution on the exterior boundary requires selecting a function that assumes $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$. Then, to determine the field equations, we utilise the function $\mathcal{F}(\mathcal{Q})$, which is expressed as follows:
$\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0\Rightarrow\mathcal{F}(\mathcal{Q})=\delta_{1}\mathcal{Q}+\delta_{2}.$ (26)
Using equations Eq. (25) and Eq. (26) with $\delta_{1}$ and $\delta_{2}$ as integration constants, the field equations for $\mathcal{F}(\mathcal{Q})$ gravity can be stated as follows:
$\rho=\frac{1}{16\pi r^{2}}\Bigg{[}2\delta_{1}+2e^{-\sigma}\delta_{1}(r{\sigma}^{\prime}-1)-r^{2}\delta_{2}\Bigg{]},$ (27)
$p_{r}=\frac{1}{16\pi r^{2}}\Bigg{[}-2\delta_{1}+2e^{-\sigma}\delta_{1}(r\Upsilon^{\prime}+1)+r^{2}\delta_{2}\Bigg{]},$ (28)
$p_{t}=\frac{e^{-\sigma}}{32\pi r}\Bigg{[}2e^{\sigma}r\delta_{2}+\delta_{1}(2+r\Upsilon^{\prime})(\Upsilon^{\prime}-\sigma^{\prime})+2r\delta_{1}\Upsilon^{\prime\prime}\Bigg{]}.$ (29)
In Eqs. (2729) prime ()($\prime$)( ′ ) denotes the differentiation with respect to radial coordinate r and $\rho$, $p_{r}$, $p_{t}$, $\sigma$ and $\Upsilon$ are unknowns which we are going to solve to get our desired result. For static and spherically symmetric fluid configurations, the general expression of the anisotropic factor, denoted as $\Delta=(p_{t}-p_{r})$, can be stated as follows:
$\Delta(r)=\frac{\delta_{1}}{8\pi}\Bigg{[}e^{-\sigma}\Bigg{(}\frac{\Upsilon^{\prime\prime}}{2}-\frac{\sigma^{\prime}\Upsilon^{\prime}}{4}+\frac{(\Upsilon^{\prime})^{2}}{4}-\frac{\Upsilon^{\prime}+\sigma^{\prime}}{2r}-\frac{1}{r^{2}}\Bigg{)}+\frac{1}{r^{2}}\Bigg{]}.$ (30)
II.2 Karmarkar Condition
The spherically symmetric line element (18) can always be embedded in six-dimensional flat spacetime, signifying its classification as belonging to embedding class two in general. On the contrary, the spherical line element can be included in five-dimensional flat spacetime, provided it fulfils the Karmarkar criterion  karmarkar1948gravitational . Then It indicates the spacetime of embedding class one. It is a necessary and sufficient condition for spherically symmetric spacetime to belong to class one. The Karmarkar condition is expressed through the curvature components as follows:
$R_{1414}=\frac{R_{1212}R_{3434}-R_{1224}R_{1334}}{R_{2323}},$ (31)
with $R_{2323}\neq 0$  pandey1982insufficiency . The nonzero components of the Riemann curvature tensor $R_{hijk}$ for the metric (18) are defined as follows:
$R_{2323}=\frac{\sin^{2}\theta(e^{\sigma}-1)r^{2}}{e^{-\sigma}}$, $R_{1212}=\frac{\sigma^{\prime}r}{2}$,
$R_{2424}=-\frac{\Upsilon^{\prime}re^{\Upsilon-\sigma}}{4}$, $R_{1334}=R_{1224}sin^{2}\theta=0$,
$R_{1414}=-\frac{1}{4e^{\Upsilon}}[2\Upsilon^{\prime\prime}+(\Upsilon^{\prime})^{2}-\sigma^{\prime}\Upsilon^{\prime}]$, $R_{3434}=\sin^{2}\theta R_{2424}.$
By substituting these components of $R_{hijk}$ in Eq. (31), we obtain the subsequent differential equation:
$\frac{2\Upsilon^{\prime\prime}+(\Upsilon^{\prime})^{2}}{\Upsilon^{\prime}}=\frac{\sigma^{\prime}e^{\sigma}}{(e^{\sigma}-1)}.$ (32)
When solving the differential Eq. (32), we get the metric potential,
$e^{\Upsilon}=\Bigg{[}A_{1}+B_{1}\int\sqrt{(e^{\sigma(r)}-1)dr}\Bigg{]}^{2},$ (33)
where $A_{1}$ and $B_{1}$ are nonzero arbitrary constants of integration. Additionally, we choose the metric potential $e^{\sigma}$ as
$e^{\sigma}=b^{2}r^{2}\text{csch}\left(fr^{2}+h\right)\text{csch}\left(fr^{2}+h\right)+1,$ (34)
where $b\neq 0$, $f\neq 0$ or $h\neq 0$  Prasad:2019epe . If b = f = h = 0, then the reduced spacetime does not constitute a class one spacetime pandey1982insufficiency . In 2017, Bhar proposed that for all well-behaved models, the metric function $e^{\sigma}$ must be monotonically rising and fulfil the constraint $e^{\sigma}(0)=1$ and $((e^{\sigma})^{\prime})_{r}=0.$ It is noted that our metric function $e^{\sigma}$ is an increasing function and satisfy the aforementioned conditions. This indicates that $e^{\sigma}$, as defined by Eq. (34), is physically acceptable. Substituting the value of $e^{\sigma}$ from Eq. (34) in Eq. (33) we ge the value of $e^{\Upsilon}$ as
$e^{\Upsilon}=\Bigg{[}A_{1}+\frac{B_{1}b}{2f}\log\Bigg{(}\frac{e^{(fr^{2}+h)}-1}{e^{(fr^{2}+h)}+1}\Bigg{)}\Bigg{]}^{2}.$ (35)
Using the value of $e^{\sigma}$ and $e^{\Upsilon}$ from Eq. (34) and Eq. (35) in Eqs. (2730), we get the expressions for energy density ($\rho$), radial pressure ($p_{r}$), tangential pressure ($p_{t}$), and the anisotropic factor ($\Delta$) as follows:
$8\pi\rho=-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\left(b^{2}r^{2}\text{csch}^{2}\left(fr^{2}+h\right)+1\right)^{2}},$
$8\pi p_{r}=\frac{8B_{1}b\delta_{1}fr^{2}\log(e)e^{fr^{2}+h}+\delta_{1}\phi_{2}}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}+\frac{\delta_{2}}{2}-\frac{\delta_{1}}{r^{2}},$
$8\pi p_{t}=\frac{\delta_{2}\phi_{5}\left(b^{4}r^{4}\text{csch}^{4}\left(fr^{2}+h\right)+1\right)+2b^{2}\phi_{6}+\phi_{4}}{2\phi_{7}\left(e^{fr^{2}+h}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)^{2}},$
$8\pi\Delta=\frac{\delta_{2}\phi_{5}\left(b^{4}r^{4}\text{csch}^{4}\left(fr^{2}+h\right)+1\right)+2b^{2}\phi_{6}+\phi_{4}}{2\phi_{7}\left(e^{fr^{2}+u}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)^{2}}$
$-\frac{8B_{1}b\delta_{1}fr^{2}\log(e)e^{fr^{2}+h}+\delta_{1}\phi_{2}}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}-\frac{\delta_{2}}{2}+\frac{\delta_{1}}{r^{2}}.$
The value of $\phi_{1}$,$\phi_{2}$,$\phi_{3}$,$\phi_{4}$,$\phi_{5}$,$\phi_{6}$ and $\phi_{7}$ are given in Appendix.
Compact star      $M(M_{0})$      $R(km)$      $\delta_{1}$      $b(km^{-1})$      $f(km^{-1})$       $h$      $\frac{M}{R}$      $A_{1}$      $B_{1}$
Cen X-3      1.48      11.6      0.2      0.107      -0.0011      1.39      0.128291379      0.0269800      0.0263286
Cen X-3      1.49      11.6      0.4      0.107      -0.0011      1.39      0.128481897      0.0235791      0.0264463
Cen X-3      1.49      11.6      0.6      0.107      -0.0011      1.39      0.128544828      0.0224455      0.0264855
Cen X-3      1.49      11.6      0.8      0.107      -0.0011      1.39      0.128576724      0.0218787      0.0265051
Cen X-3      1.49      11.6      1.0      0.107      -0.0011      1.39      0.128671552      0.0201782      0.026564
Table 1: The numerical value of Mass $M(M_{0})$, Radius $R(km)$, $b$, $f$, $h$, $A_{1}$ and $B_{1}$ for different values of $\delta_{1}$
Expression Remarks
$\rho,p_{r},p_{t}$ $>0$, acceptable
$\frac{p_{rc}}{\rho_{c}}$ $\leq 0$
$\Delta$ $>0$, acceptable
$\frac{d\rho}{dr},\frac{dp_{r}}{dr},\frac{dp_{t}}{dr}$ $<0$, acceptable
$F_{a},F_{h},F_{g}$ balanced
$\rho\pm p_{r}$ $>0$, acceptable
$\rho\pm p_{t}$ $>0$, acceptable
$\rho+p_{r}+2p_{t}$ $>0$, acceptable
$m(r)$ $>0,acceptable$
$u(r)$              $0<u(r)<\frac{8}{9}$, acceptable
$Z_{S}$            $50<Z_{S}<$ 5, acceptable
$\omega_{r},\omega_{t}$               $10<\omega_{r},\omega_{t}<$ 1, acceptable
$V^{2}_{r},V^{2}_{t}$                 $10<V^{2}_{r},V^{2}_{t}<$ 1, acceptable
$V^{2}_{t}-V^{2}_{r}$                  $-1<|V^{2}_{t}-V^{2}_{r}|<1$, acceptable
$\Gamma$ $>\frac{4}{3}$, acceptable
Table 2: The table shows the acceptability of all parameters of the proposed compact star
III Physical properties and characteristics of the model
In this section, we describe the physical properties and characteristics of an anisotropic compact star.
III.1 Metric elements
Anisotropic compact stars require finite and nonsingular metric elements to produce a physically plausible model. In our model, the value of $e^{\sigma}$ at $r=0$ equals 1, and the value of $e^{\Upsilon}$ at r=0 equals 0.4583. we have a positive value of $e^{\sigma}$ and $e^{\Upsilon}$ at r=0, indicating that the metric elements are stable at the centre of the compact star. The Exact behaviour of the metric elements has been shown in Fig. 1. The graph of metric elements $e^{-\sigma}$ and $e^{\Upsilon}$ are finite and nonsingular throughout the radius of the compact star. The graph of $e^{-\sigma}$ is monotonically decreasing towards the surface and the graph of $e^{\Upsilon}$ monotonically increases towards the surface. Both the graph coincides at the boundary of the compact star from these characteristics of metric elements we can conclude that metric elements are competent to produce the model for the anisotropic compact star.
Figure 1: Graphical representation of metric elements $e^{-\sigma}$ and $e^{\Upsilon}$ for fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and radius r of the proposed compact star.
III.2 Seamless Matching of Interior and Exterior spacetime
The boundary requirement for solving the Einstein-field equations in $\mathcal{F}(\mathcal{Q})$ gravity is presented in this section. To accomplish this, we compare the solution on the exterior and interior spacetimes. There are numerous indications in more recent works, such as Wang:2021zaz , that the feasible solution in $\mathcal{F}(\mathcal{Q})$ gravity is solvable using Schwarzchild (Anti-) di sitter solution. The solution is provided as follows:
$ds^{2}=-\Bigg{(}1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}\Bigg{)}dt^{2}+\frac{dr^{2}}{\Bigg{(}1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}\Bigg{)}}$ (40)
$+\hskip 2.84544ptr^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$
Taking $r=R$, the Schwarzchild (Anti-) di sitter metric has been solved at the surface using the requirement that Darmois-Israel derived in their research  Israel:1966rt
$\Bigg{(}1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}\Bigg{)}=e^{\Upsilon(R)},$ (41)
$\Bigg{(}1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}\Bigg{)}=e^{-\sigma(R)},$ (42)
$\hskip 42.67912ptp_{r}(R)=0.$ (43)
To ensure continuity of the first derivative:
$\Bigg{(}\frac{\partial w^{-}_{ss}}{\partial r}\Bigg{)}_{r=R}=\quad\Bigg{(}\frac{\partial w^{+}_{ss}}{\partial r}\Bigg{)}_{r=R}$ (44)
and
$w^{-}_{ss}=e^{\Upsilon},\quad w^{+}_{ss}=\Bigg{(}1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}\Bigg{)}.$ (45)
Now from Eq. (44) and Eq. (45) we get,
$\Upsilon^{\prime}(R)=\frac{6M-2\Lambda R^{3}}{3R^{2}-6MR-\Lambda R^{4}}$ (46)
By solving equation Eq. (28) and Eq. (43) weget,
$\Upsilon^{\prime}(R)=\frac{6M-2\Lambda R^{3}}{3R^{2}-6MR-\Lambda R^{4}}$ (47)
The analysis from Eq. (46) and Eq. (47) demonstrates the continuity of the first derivative at $r=R.$
Where the notations $M$ and $\Lambda$ denote the mass and cosmological constant.
It is noteworthy that the cosmological constant $\Lambda$ is contingent upon the constants $\delta_{1}$ and $\delta_{2}$, which are expressed as follows $\Lambda=\frac{\delta_{2}}{2\delta_{1}}$. In this scenario, the evidence for the cosmological constant is zero, hence its value does not affect the current stellar model. Normally, the cosmological constant would have a significant value to influence the mathematical formulation of the problem. The cosmological constant, which is zero in the current study, is estimated to be $10^{-46}km^{-2}$, based on numerical evidence. For this project, $\delta_{2}$ has been set to 0.00001 for the entire problem. $\delta_{1}$ has a significant impact on this proposed model because as we increase the value of $\delta_{1}$, the mathematical values and graphical behaviours of physical properties like density, pressure, adiabatic index, mass, compactness, and velocity of sound also increase. Along with this, the mathematical values of central density and central pressure of the compact star also increase.
Eqs. (4143) are boundary conditions with the help of these boundary conditions we calculated the values of the constants $A_{1}$ and $B_{1}$ that is
$A_{1}=\frac{B_{1}b\left(-\phi_{10}\left(\delta_{2}+\phi_{11}\right)-16\delta_{1}f\log(e)e^{fR^{2}+h}\right)}{2f\phi_{12}\left(e^{2\left(fR^{2}+h\right)}-1\right)},$ (49)
$B_{1}=-\frac{\phi_{13}e^{-fR^{2}-h}\left(e^{2\left(fR^{2}+h\right)}-1\right)}{8b\delta_{1}\log(e)}.$
The value of $\phi_{10}$,$\phi_{11}$,$\phi_{12}$, and $\phi_{13}$ are given in appendix
III.3 Behaviour of density
For a well-behaved model of an anisotropic compact star, the density value is maximum at the centre and decreases as the radius increases towards the surface. In our model, the density value is Optimum at the centre and decreases as the radius increases towards the surface as shown in Fig. 2. Therefore our model is well-behaved.
III.4 Behaviour of radial and tangential pressure
. Here we analyse the graphical behaviour of radial pressure ($p_{r}$) and tangential pressure ($p_{t}$). Figure 2 represents the behaviour of radial and tangential pressure, both pressures attain the Optimum value at the centre(r=0) and decrease monotonically as the radius increases towards the surface. Tangential pressure is always positive throughout the compact star, but radial pressure diminishes toward the border.
III.5 Anisotropy factor and anisotropic force
In a static and spherically symmetric non-perfect fluid structure, the anisotropy factor, $\Delta=(p_{t}-p_{r})$, defines the dissimilarity in pressures, particularly the radial pressure ($p_{r}$) and the tangential pressure ($p_{t}$).
A force due to the anisotropic pressure is represented by $\Delta/r$, which is repulsive, if $p_{t}>p_{r}\leftrightarrow\Delta>0$, and attractive if $p_{t}<p_{r}\leftrightarrow\Delta<0$ of the stellar model. For the considered matter distribution when $p_{t}>p_{r}$ allows the construction of more compact objects, compared to isotropic fluid sphere Gokhroo:1994fbj ; Maurya:2017uyk ; maurya2017anisotropic ; Maurya:2017jdo ; morales2018compact etc. A positive anisotropy factor improves the stability and equilibrium of the system. The value of $\Delta$ must be zero in the centre of the star structure. Figure 2 indicates that the value of the anisotropy factor $\Delta$ is zero at the centre of the compact star and the value of the anisotropy factor $\Delta$ is positively rising within the star interior of our model, suggesting that the system undergoes a repulsive anisotropic force.
Figure 2: The behaviour of energy density($\rho$), radial pressure($p_{r}$), tangential pressure($p_{t}$) and anisotropy($\Delta$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
Figure 3: The behaviour of gradient of energy density($\frac{d\rho}{dr}$), gradient of radial pressure($\frac{dp_{r}}{dr}$) and gradient of tangential pressure($\frac{dp_{t}}{dr}$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
III.6 Zel’dovich requirement
The Zel‘dovich ratio is calculated for both radial and tangential orientations  Zeldovich:1968ehl . The model of the compact star evidently fulfils these two requirements, as shown in Fig. 4. Regarding the interior of stars, where $r<R$, we utilise $\omega_{r}=\frac{p_{r}}{\rho}<1$ and $\omega_{t}=\frac{p_{t}}{\rho}<1$, both of which are less than 1. Consequently, the model of a compact star can be sustained against gravitational collapse with support from degeneracy. The expression derived numerically for $\omega_{r}$ and $\omega_{t}$ are:
$\omega_{r}=\Bigg{(}\frac{8B_{1}b\delta_{1}fr^{2}\log(e)e^{fr^{2}+h}+\delta_{1}\phi_{2}}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}+A\Bigg{)}B,$ (51)
$\omega_{t}=\Bigg{(}\frac{\delta_{2}\phi_{5}\left(b^{4}r^{4}\text{csch}^{4}\left(fr^{2}+h\right)+1\right)+C}{2\phi_{7}\left(e^{fr^{2}+h}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)^{2}}\Bigg{)}B.$
Figure 4: The behaviour $\omega_{r}$ and $\omega_{t}$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
III.7 Energy bounds
When analysing data, all energy constraints must be met for the model to be feasible. Here NEC represents the null energy condition, WEC represents the weak energy condition, DEC represents the dominant energy condition, SEC represents the strong energy condition and TEC depicts the trace energy condition.
Figure 5: The behaviour of NEC, WEC, DEC, SEC and TEC for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
$NEC:\rho\geq 0,$ (52)
$WEC:\rho+p_{r}\geq 0,\rho+p_{t}\geq 0,$ (53)
$DEC:\rho-p_{r}\geq 0,\rho-p_{t}\geq 0,$ (54)
$SEC:\rho+p_{r}+2p_{t}\geq 0,$ (55)
$TEC:\rho-p_{r}-2p_{t}\geq 0.$ (56)
Figure 5 represents the energy requirements graphically; all energy requirements needs have been identified and are fully met. Therefore, our solutions are physically possible.
The mathematical expression for $\rho$ is given in Eq. (II.2), while the expressions for $\rho+p_{r}$, $\rho+p_{t}$, $\rho-p_{r}$, $\rho-p_{t}$, $\rho+p_{r}+2p_{t}$, and $\rho-p_{r}-2p_{t}$ are provided in the appendix.
III.8 Adiabatic index
The adiabatic index is an essential parameter that defines the stability of a stellar structure. For a spherically symmetric object, testing the adiabatic index is crucial since it indicates its solidity. For a stable compact star structure the value of the adiabatic index( $\Gamma$) should be greater than $\frac{4}{3}$ to guarantee stability heintzmann1975neutron .The adiabatic index is often defined as:
$\Gamma<\frac{4}{3}-\Bigg{[}\frac{4}{3}\frac{p_{r}-p_{t}}{|p_{r}^{\prime}|r}\Bigg{]}_{max}$ (57)
Our present investigation examines the graphical behaviour of the adiabatic index, shown in Fig. 6. The value of the adiabatic index($\Gamma$) is greater than $\frac{4}{3}$ for any value of r. From this, we can say that our proposed model satisfies the required stability criterion($\Gamma>\frac{4}{3}$). The mathematical expression for the adiabatic index($\Gamma$) is given in the appendix.
Figure 6: Graphical representation of Adiabatic index $\Gamma$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
III.9 Stability and Equilibrium
This subsection discusses the graphical depiction of stability and equilibrium in the stellar model utilising TOV equations, causality Condition, Stability Versus Convection and Harrison-Zeldovich-Novikov Condition in the interior of a compact star.
III.9.1 Equilibrium requirement
The Tolman-Oppenheimer-Volkoff (TOV) equation has been described as an equilibrium condition for a compact star, combining gravitational, hydrostatic and anisotropic forces  zubair2016possible . The standard representation of the TOV equation is expressed as
$\frac{dp_{r}}{dr}+\frac{\Upsilon^{\prime}(\rho+p_{r})}{2}-\frac{2(p_{t}-p_{r})}{r}=0$ (58)
On the other hand, it can be stated as
$F_{g}+F_{h}+F_{a}=0$; $F_{g}=-\frac{\Upsilon^{\prime}(\rho+p_{r})}{2}$, $F_{h}=-\frac{dp_{r}}{dr}$, $F_{a}=\frac{2(p_{t}-p_{r})}{r}.$
This $F_{g}$, $F_{h}$ and $F_{a}$ indicate the gravitational, hydrostatic, and anisotropic forces, respectively. Using the value of $\rho$, $p_{r}$ and $p_{t}$ in the TOV equation, we represent these forces visually shown in Fig. 7. Each of these forces has small positive or negative values; nevertheless, together, they satisfy the TOV equation. Therefore, we find that the revealed stellar structure is in a condition of equilibrium as a result of these forces.
The mathematical expression for $F_{g}$, $F_{h}$ and $F_{a}$ are given in the appendix.
Figure 7: The behaviour of gravitational force ($F_{g}$), hydrostatic force ($F_{h}$) and anisotropic force ($F_{a}$) versus the radius r of the proposed compact star and $\delta_{1}$ for a fixed value of $\delta_{2}=0.00001$.
III.9.2 Causality and Stability requirement
This study examines the validity of the causality criterion for the proposed stellar model. The causality conditions provide boundaries on the radial and tangential velocities, denoted as $V^{2}_{r}$ and $V^{2}_{t}$, respectively. The boundaries are specified as $10<|V^{2}_{i}|<$ 1, where $i=r,t$. The $V^{2}_{r}$ and $V^{2}_{t}$ are defined by
$V^{2}_{r}=\frac{dp_{r}}{d\rho},\quad V^{2}_{t}=\frac{dp_{t}}{d\rho}$ (59)
The graphical representation of these velocities is illustrated in Fig. 8, demonstrating the validity of causality criteria. Furthermore,  Abreu:2007ew has proposed an additional stability requirement for these velocities, denoted as $-1\leq V^{2}_{t}-V^{2}_{r}\leq 1$. This behaviour is regarded as a significant characteristic of compact star models. In the present case, we visually examine this state as shown in Fig. 8. The graph strongly illustrates that our model fulfils this stability criteria as well.
The mathematical expression for $V^{2}_{r}=\frac{dp_{r}}{d\rho}$ and $V^{2}_{t}=\frac{dp_{r}}{d\rho}$ are given in the appendix.
Figure 8: The behaviour of radial velocity ($V^{2}_{r}$), Tangential velocity ($V^{2}_{t}$ ) and Resultant velocity($|V^{2}_{t}-V^{2}_{r}|$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
III.9.3 Stability Versus Convection
The buoyancy principle inside a fluid demonstrates that every displaced fluid component will revert to its original position, as evidenced by the stability of a self-gravitating sphere in contrast to convection. It was illustrated in  Hernandez:2018sjc , such that
$\rho^{\prime\prime}(r)\leq 0$ (60)
We demonstrate the graphical representation of $\rho^{\prime\prime}$ as a function of r in Fig. 9. The model demonstrates stability in the inner shells during convective motion, whereas the outside shells remain unstable  Andrade:2022idg .
The mathematical expression for $\rho^{\prime\prime}$ is given in the appendix.
Figure 9: The behaviour of $\rho^{\prime\prime}$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
III.9.4 Harrison-Zeldovich-Novikov Stability Condition
Harrison  harrison1965gravitation and Zeldovich-Novikov  zeldovich1971relativistic proposed stability criteria for the compact star model, contingent upon the star’s mass and centre density. According to these observations, they suggested that for a stable configuration, $\frac{dM}{d\rho_{c}}>0$, where M and $\rho_{c}$ represent the mass and central density of the compact star, respectively. With the central density, we will define the mass function of our static anisotropic fluid solution as
$M(\rho_{c})=\frac{R}{2}\left(-\frac{D}{E\left(2\rho_{c}+\delta_{2}\right)+D}-\frac{\delta_{2}R^{2}}{6\delta_{1}}+1\right).$ (61)
When differentiating Eq. (61) with respect to $\rho_{c}$, we obtain
$\frac{dM}{d\rho_{c}}=\frac{4EDR}{\left(4E\rho_{c}+2E\delta_{2}+12\delta_{1}\text{csch}^{2}(h)\right){}^{2}}.$ (62)
Figure 10 demonstrates that our anisotropic models fulfil the Harrison-Zeldovich-Novikov stability criterion. The stability of our structures increases with increasing radii and remains constant after the central densities attain their maximum values.
Figure 10: The behaviour of Mass and $\frac{dM}{d\rho_{c}}$ for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the central density $\rho_{c}$ of the proposed compact star.
Compact star   $\delta_{1}$    density value at r=0($gm/cm^{3}$)       density value at r=R($gm/cm^{3}$)   pressure value at r=0($dyne/cm^{2}$)
Cen X-3   0.2    0.041349 ×$\times 10^{14}$      0.03222×$\times 10^{14}$   0.043506×$\times 10^{34}$
Cen X-3   0.4   0.082698×$\times 10^{14}$      0.06444×$\times 10^{14}$   0.091846×$\times 10^{34}$
Cen X-3   0.6   0.124047×$\times 10^{14}$       0.09666 ×$\times 10^{14}$    0.140186 ×$\times 10^{34}$
Cen X-3   0.8   0.165396×$\times 10^{14}$      0.12888×$\times 10^{14}$   0.188526×$\times 10^{34}$
Cen X-3   1.0   0.206745×$\times 10^{14}$      0.16110×$\times 10^{14}$   0.236866×$\times 10^{34}$
Table 3: The value of central density($\rho_{c}$), surface density and the central pressure($p_{rc}$) for different values of $\delta_{1}$
III.10 Redshift Profile
A phenomenon known as gravitational redshift occurs when electromagnetic waves or photons, lose energy when they emerge from a gravitational well. Electromagnetic radiation expands its wavelength and photon energy to escape when it moves out of a gravitational well. It dissipates energy not by altering its velocity but by varying its frequency, as it must propagate concurrently at the speed of light. When energy is diminished, the photon’s wave frequency declines and its wavelength elongates, leading to a shift toward the red end of the electromagnetic spectrum, referred to as redshift. The behaviour of interior redshift similarly reflects the profile of interior density. A photon must traverse a longer distance and pass through a denser part of the compact star’s interior to exit the centre and reach the surface. This greater dispersion causes a significant energy loss. A photon that ascends near the surface will traverse a shorter distance through a less dense region, leading to less dispersion and energy loss. This means that the interior redshift is highest in the centre and lowest toward the surface. The surface redshift depends on the surface gravity, which is defined by the stellar object’s radius and total mass. Surface gravity increases as a result of a modest increase in radius, similar to the increase in mass. As a result, the surface redshift increases monotonically towards the surface of the star. Gravitational and surface redshifts are defined as:
$Z_{G}=\frac{1}{\sqrt{|e^{\Upsilon(r)}|}}-1=\frac{1}{A_{1}+\frac{B_{1}b}{2f}\log\Bigg{(}\frac{e^{(fr^{2}+h)}-1}{e^{(fr^{2}+h)}+1}\Bigg{)}}-1,$ (63)
$Z_{S}=\frac{1}{\sqrt{1-2u(r)}}-1.$ (64)
Where the value of $u(r)$ is given in Eq. (68), the graph of gravitational redshift is monotonically decreasing towards the surface and the graph of surface redshift is monotonically increasing towards the surface as shown in Fig. 11.
Figure 11: The behaviour of Gravitational Redshift ($Z_{G}$) and Surface Redshift ($Z_{S}$) for a fixed value of $\delta_{2}=0.00001$ versus $\delta_{1}$ and the radius r of the proposed compact star.
III.11 Mass and compactness of the model
The compact star’s compactness and mass are shown in Fig. 12. The impact of parameter $\delta_{1}$ on the compactness and mass of the compact star is visible. When the value of $\delta_{1}$ increases, mass and compactness also increase. The mass and compactness values are zero in the centre of the compact star and rise monotonically toward the surface. It is significant to note that there is a mass-radius ratio bound for any stellar structure, which is provided by Buchdahl and is $\frac{2M}{R}\leqslant\frac{8}{9}$ Buchdahl:1959zz This bound has been computed for the compact star in our model and is displayed in Table 1. This bound is less than $\frac{4}{9}$, demonstrating that our model satisfies Buchdahl’s requirements. The functions $m(r)$, $u(r)$ of radial coordinate r are used to represent the compact star mass and compactness are given below:
[…]
The following is the expression for the mass and compactness of the compact star:
$m(r)=-\frac{2\pi r^{3}\left(2b^{2}\left(\delta_{2}r^{2}-6\delta_{1}\right)+\delta_{2}\left(\phi_{14}-1\right)\right)}{3\left(2b^{2}r^{2}+\phi_{14}-1\right)},$ (67)
$u(r)=-\frac{2\pi r^{2}\left(2b^{2}\left(\delta_{2}r^{2}-6\delta_{1}\right)+\delta_{2}\left(\phi_{14}-1\right)\right)}{3\left(2b^{2}r^{2}+\phi_{14}-1\right)}.$ (68)
The star obtained here is inside the star’s limit of being a compact star, with a solar mass of 1.49 for a value of $\delta_{1}$=1. The graph’s growing tendency suggests that the star’s mass rises gradually as the parameter $\delta_{1}$ rises. In this scenario, the constant $\delta_{2}$ is supposed to be 0.00001, therefore it has minimal impact on the mass and compactness of the model.
Figure 12: The behaviour of mass function $m(r)$ and compactness parameter $u(r)$ versus the radius r of the proposed compact star and $\delta_{1}$ for a fixed value of $\delta_{2}=0.00001$.
IV Conclusion
This study examines a spherically symmetric compact star model with non-perfect fluid. We examined the generic function $\mathcal{F}(\mathcal{Q})$ defined as $\mathcal{F}(\mathcal{Q})=\delta_{1}\mathcal{Q}+\delta_{2}$, where $\delta_{1}$ and $\delta_{2}$ are integration constant. The Karmarkar criteria have been used to develop the differential equation connecting the metric elements of spacetime. By solving this differential equation, we have derived an expression for the metric element $e^{\Upsilon}$ in relation to $e^{\sigma}$. The juxtaposition of the internal spherical metric with the exterior Schwarzschild metric imposed constraints on the arbitrary constants throughout this structure. The results acquired were illustrated visually in Fig. 1– 12. The main results may be described as follows:
  • Figure 1 illustrates that the graph of $e^{-\sigma}$ is regular, finite, and decreasing towards the surface, whereas the graph of $e^{\Upsilon}$ is regular, finite, and growing towards the surface, with both graphs coinciding at the boundary.
  • The energy density and both radial and tangential pressure of the star are positive, finite, and monotonically decreasing throughout its entirety, as seen in Fig. 2. The radial pressure diminishes toward the star’s boundary, with the central pressure approximating $10^{34}dyne/cm^{2}$ and the central density about $10^{14}g/cm^{3}$, as seen in Table 3.
  • We observe from Fig. 2 that the anisotropy factor $\Delta$ vanishes at the centre of the star and remains positive throughout the star, showing the presence of opposing forces. The graphical behaviour of the gradients of density, radial pressure, and tangential pressure is illustrated in Fig. 3.
  • The value of $\omega_{r}$ and $\omega_{t}$ are greater than zero and less than one that is $10<\omega_{r},\omega_{t}<$ 1 and both the graph are monotonically decreasing towards the surface as shown in Fig. 4.
  • The energy criteria, including the null energy condition (NEC), weak energy condition (WEC), dominant energy condition (DEC), strong energy condition (SEC) and trace energy condition (TEC) are fulfilled for our anisotropic models are shown in Fig. 5.
  • The adiabatic index $\Gamma$ for our model exceeds 4/3 at all interior points within the star, shown in Fig. 6 , therefore it verify the stability of our model.
  • We analyze the TOV equation to characterise the equilibrium condition influenced by gravitational force $(F_{g})$, hydrostatic force $(F_{h})$ and anisotropic forces $(F_{a})$. Figure 7 illustrates that the $(F_{g})$ is counterbalanced by the combined effects of $(F_{h})$ and $(F_{a})$ to achieve the necessary stability of the model. Nevertheless, the impact of $(F_{a})$ is less than that of $(F_{h})$.
  • The radial and tangential sound velocities in our model obey the causality criterion, namely $V^{2}_{r},V^{2}_{t}<1.$ They also obey the Abreu constraint: $-1\leq V^{2}_{t}-V^{2}_{r}\leq 1$. The graph of $V^{2}_{r},V^{2}_{t}$ are monotonically decreasing towards the surface as shown in Fig. 8. For a stable compact star, the condition for stability against convection requires that the stability criterion be negative throughout the stellar interior. In our model, the value of the stability criterion is negative throughout the star, as shown in Fig. 9, indicating that the model is stable against convective instability.
  • We observe from Fig. 10 our model satisfy Harrison-Zeldovich-Novikov condition that is $\frac{dM}{d\rho_{c}}>0$.
  • Figure 11 illustrates that gravitational redshift exhibits a monotonically declining trend, but surface redshift demonstrates a monotonically growing trend as one approaches the surface. Gravitational redshift is greatest in the centre and smallest at the boundary of the compact star, whereas surface redshift is smallest at the centre and greatest at the boundary of the model.  Boehmer:2006ye assert that the redshift must fulfil the requirement $Z\leq 5$ and in the model of an anisotropic star, Ivanov  Ivanov:2002xf proposed that the surface redshift is limited to a maximum of 3.842 when the pressure $p_{t}$ adheres to the SEC and 5.211 when $p_{t}$ complies with the DEC. Figure 11 demonstrates that the Gravitational and Surface redshift of our models are in substantial concordance.
  • The compact star’s compactness and mass are shown in Fig. 12. The impact of parameter $\delta_{1}$ on the compactness and mass of the compact star is visible. When the value of $\delta_{1}$ increases, the value of mass and compactness also increases. The values of mass and compactness are zero at the centre of the compact star, and they monotonically increase towards the surface. It is significant to note that there is a mass-radius ratio bound for any stellar structure, which is provided by Buchdahl and is $\frac{2M}{R}\leqslant\frac{8}{9}$ Buchdahl:1959zz . This bound has been computed for the compact star in our model and is displayed in Table 1. This bound is less than $\frac{4}{9}$, demonstrating that our model complies with Buchdahl’s requirements.
  • We have examined the behaviour of various quantities, including metric elements, energy density, radial and tangential pressures, energy and pressure gradients concerning the radial coordinate, anisotropy parameter, $\omega_{r}$ and $\omega_{t}$, energy conditions, adiabatic index, TOV equation, causality conditions, Harrison-Zeldovich-Novikov condition, redshift, mass and compactness for the astronomical object Cen X-3. The results are summarized in Table 2, which confirms the physical plausibility and mathematical coherence of our spherically symmetric model.
Acknowledgments
Sat Paul acknowledges the Department of Mathematics, Central University of Haryana, India, for providing the necessary support for writing and finalising the paper. I also acknowledge the Ministry of Tribal Affairs, New Delhi, India, for providing financial support. Jitendra Kumar acknowledges the authority of the Central University of Haryana (CUH). S. K. Maurya thanks the administration of the University of Nizwa in the Sultanate of Oman for their constant support and encouragement.
Data access statement
This manuscript includes no accompanied data, therefore the data will not be deposited. The essential calculations and graphic description are already included in the paper.
Conflict of interest
In this research paper, The authors have no conflicts of interest.
Appendix
$\phi_{1}=2b^{2}\text{csch}^{2}\left(fr^{2}+h\right)\left(-3\delta_{1}+4\delta_{1}fr^{2}\coth\left(fr^{2}+h\right)+\delta_{2}r^{2}\right)$
$\phi_{2}=\left(e^{2\left(fr^{2}+h\right)}-1\right)\left(2A_{1}f+B_{1}b\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)\right)$
$\phi_{3}=\left(b^{2}r^{2}\text{csch}^{2}\left(fr^{2}+h\right)+1\right)\left(2A_{1}f+B_{1}b\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)\right)$
$\phi_{4}=16B_{1}b\delta_{1}f\log(e)e^{fr^{2}+h}\left(e^{2\left(fr^{2}+h\right)}-1\right)-16B_{1}b\delta_{1}f^{2}r^{2}\log^{2}(e)e^{fr^{2}+h}\left(e^{2\left(fr^{2}+h\right)}+1\right)$
$\phi_{5}=\left(e^{2\left(fr^{2}+h\right)}-1\right)^{2}\left(2A_{1}f+B_{1}b\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)\right)$
$\phi_{6}=\text{csch}^{2}\left(fr^{2}+h\right)\left(4B_{1}b\delta_{1}fr^{2}\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}-8B_{1}\phi_{8}+2\delta_{1}\phi_{9}-\phi_{5}\left(\delta_{1}-\delta_{2}r^{2}\right)\right)$
$\phi_{7}=\left(b^{2}r^{2}\text{csch}^{2}\left(fr^{2}+h\right)+1\right)^{2}\left(2A_{1}f+B_{1}b\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)\right)$
$\phi_{8}=b\delta_{1}f^{2}r^{4}\log^{2}(e)e^{fr^{2}+h}\left(e^{2\left(fr^{2}+h\right)}+1\right)$
$\phi_{9}=2\delta_{1}fr^{2}\left(e^{2\left(fr^{2}+h\right)}-1\right)\coth\left(fr^{2}+h\right)\left(4B_{1}bfr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$\phi_{10}=\left(e^{2\left(fR^{2}+h\right)}-1\right)\log\left(\frac{e^{fR^{2}+h}-1}{e^{fR^{2}+h}+1}\right)$
$\phi_{11}=b^{2}\left(2\delta_{1}-\delta_{2}R^{2}\right)\text{csch}^{2}\left(fR^{2}+h\right)$
$\phi_{12}=\delta_{1}-b^{2}\left(2\delta_{1}-\delta_{2}R^{2}\right)\text{csch}^{2}\left(fR^{2}+h\right)$
$\phi_{13}=\sqrt{\frac{1}{b^{2}R^{2}\text{csch}^{2}\left(fR^{2}+h\right)+1}}\left(\delta_{2}-b^{2}\left(2\delta_{1}-\delta_{2}R^{2}\right)\text{csch}^{2}\left(fR^{2}+h\right)\right)$
$\phi_{14}=\cosh\left(2\left(fr^{2}+h\right)\right)$
$A=\frac{\delta_{2}}{2}-\frac{\delta_{1}}{r^{2}}$
$B=-\frac{2\left(b^{2}r^{2}\text{csch}^{2}\left(fr^{2}+h\right)+1\right)^{2}}{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}$
$C=2b^{2}\phi_{6}+\phi_{4}$
$D=6\delta_{1}\text{csch}^{2}(h)$
$E=R^{2}\text{csch}\left(fR^{2}+h\right)\text{csch}\left(fR^{2}+h\right)$
$\phi_{15}=1+b^{2}r^{2}\text{csch}\left(fr^{2}+h\right)^{2}$
$\phi_{16}=aB_{1}\delta_{1}\left(\frac{2fr\log(e)e^{fr^{2}+h}}{e^{fr^{2}+h}+1}-\frac{2fr\log(e)e^{fr^{2}+h}\left(e^{fr^{2}+h}-1\right)}{\left(e^{fr^{2}+h}+1\right)^{2}}\right)\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$\phi_{18}=\frac{2\delta_{1}f\log(e)e^{fr^{2}+h}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)^{2}}$
$\phi_{19}=\delta_{1}\left(2b^{2}fr\text{csch}^{2}\left(fr^{2}+h\right)-4b^{2}fr^{3}\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\right)\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$\phi_{20}=A-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+\frac{\text{$\delta$1}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}$
$\phi_{21}=A+\frac{\delta_{1}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}$
$\phi_{22}=\frac{2fr\log(e)e^{fr^{2}+h}}{e^{fr^{2}+h}+1}-\frac{2fr\log(e)e^{fr^{2}+h}\left(e^{fr^{2}+h}-1\right)}{\left(e^{fr^{2}+h}+1\right)^{2}}$
$\phi_{23}=\frac{bB_{1}\phi_{22}\left(e^{fr^{2}+h}+1\right)\left(e^{2\left(fr^{2}+h\right)}-1\right)}{e^{fr^{2}+h}-1}$
$\phi_{24}=\frac{2\delta_{1}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{3}\phi_{3}\left(e^{br^{2}+h}+1\right)\left(e^{fr^{2}+h}-1\right)}$
$\phi_{25}=\frac{4fr\log(e)e^{2\left(fr^{2}+h\right)}\left(2A_{1}f+bB_{1}\log\left(\frac{e^{br^{2}+c}-1}{e^{br^{2}+c}+1}\right)\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}$
$\phi_{26}=b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+2b^{2}\text{csch}^{2}\left(fr^{2}+h\right)\left(-3\delta_{1}+4\delta_{1}fr^{2}\coth\left(fr^{2}+h\right)+\delta_{2}r^{2}\right)+\delta_{2}$
$\phi_{27}=2b^{4}r\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+2b^{4}\delta_{2}r^{3}\text{csch}^{4}\left(fr^{2}+h\right)$
$-8b^{4}fr^{3}\left(\delta_{2}r^{2}-2\delta_{1}\right)\coth\left(fr^{2}+h\right)\text{csch}^{4}\left(fr^{2}+h\right)$
$+2b^{2}\text{csch}^{2}\left(fr^{2}+h\right)\left(-8\delta_{1}f^{2}r^{3}\text{csch}^{2}\left(fr^{2}+h\right)+8\delta_{1}fr\coth\left(fr^{2}+h\right)+2\delta_{2}r\right)$
$-8b^{2}fr\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(-3\delta_{1}+4\delta_{1}fr^{2}\coth\left(fr^{2}+h\right)+\delta_{2}r^{2}\right)$
$\phi_{28}=2b^{2}r\text{csch}^{2}\left(fr^{2}+h\right)-4b^{2}fr^{3}\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)$
$\phi_{29}=-\frac{\phi_{19}}{r^{2}\phi_{7}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}$
$\phi_{30}=\frac{2\delta_{1}}{r^{3}}-\frac{\phi_{16}}{r^{2}\phi_{5}\phi_{15}}-2\phi_{18}-\phi_{24}-\phi_{29}$
$\Gamma=\frac{\phi_{20}\phi_{28}\phi_{26}\left(\delta_{1}\left(16bB_{1}f^{2}r^{3}\log^{2}(e)e^{fr^{2}+h}+16bB_{1}fr\log(e)e^{fr^{2}+h}+\phi_{23}+\phi_{25}\right)+\phi_{30}\right)}{\phi_{21}\phi_{15}^{3}}-\frac{\phi_{27}}{2\phi_{15}^{2}}$
$\rho+p_{r}=A-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+\frac{\delta_{1}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}$
$W_{1}=b^{4}\delta_{2}r^{4}\phi_{5}\text{csch}^{4}\left(fr^{2}+h\right)-16bB_{1}\delta_{1}f^{2}r^{2}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}$
$+16bB_{1}\delta_{1}f\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}+\delta_{2}\phi_{5}$
$W_{2}=-8bB_{1}\delta_{1}f^{2}r^{4}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}$
$+2\delta_{1}fr^{2}\left(e^{2\left(fr^{2}+h\right)}-1\right)\coth\left(fr^{2}+h\right)\left(4bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$+4bB_{1}\delta_{1}fr^{2}\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}-\phi_{5}\left(\delta_{1}-\delta_{2}r^{2}\right)$
$\rho+p_{t}=-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(br^{2}+c\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+\frac{2b^{2}W_{2}\text{csch}^{2}\left(fr^{2}+h\right)}{2\phi_{7}\left(e^{br^{2}+c}-1\right)^{2}\left(e^{br^{2}+c}+1\right)^{2}}+W_{1}$
$\rho+p_{r}+2p_{t}=-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+\frac{2b^{2}W_{2}\text{csch}^{2}\left(fr^{2}+h\right)}{\phi_{7}\left(e^{fr^{2}+h}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)^{2}}$
$+\frac{\delta_{1}\left(8bf\text{B1}r^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}+A+W_{1}$
$\rho-p_{r}=-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+A-\frac{\delta_{1}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}$
$\rho-p_{t}=-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+A+\frac{2b^{2}W_{2}\text{csch}^{2}\left(fr^{2}+h\right)}{\phi_{7}\left(e^{fr^{2}+h}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)^{2}}$
$+\frac{\delta_{1}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}+W_{1}$
$W_{3}=b^{4}\delta_{2}r^{4}\phi_{5}\text{csch}^{4}\left(fr^{2}+h\right)-16bB_{1}\delta_{1}f^{2}r^{2}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}$
$-16bB_{1}\delta_{1}f\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}+\delta_{2}\phi_{5}$
$\rho-p_{r}-2p_{t}=A-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+b^{4}\delta_{2}r^{4}\phi_{5}\text{csch}^{4}\left(fr^{2}+h\right)$
$+\frac{2b^{2}W_{2}\text{csch}^{2}\left(fr^{2}+h\right)}{\phi_{7}\left(e^{fr^{2}+h}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)^{2}}-\frac{\text{$\delta$1}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}-W_{3}$
$Z_{1}=A-\frac{b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}}{2\phi_{15}^{2}}+\frac{\phi_{24}}{2}$
$Z_{2}=\frac{2\delta_{1}f\log(e)e^{fr^{2}+h}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r\phi_{3}\left(e^{fr^{2}+h}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)}$
$Z_{3}=16bB_{1}f^{2}r^{3}\log^{2}(e)e^{fr^{2}+h}+16bB_{1}fr\log(e)e^{fr^{2}+h}$
$Z_{4}=\left(A_{1}+\frac{bB_{1}\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)}{2f}\right)$
$Z_{5}=\frac{bB_{1}\phi_{22}\left(e^{fr^{2}+h}+1\right)\left(e^{2\left(fr^{2}+h\right)}-1\right)}{e^{fr^{2}+h}-1}$
$Z_{6}=2b^{4}\delta_{2}r^{3}\text{csch}^{4}\left(fr^{2}+h\right)-8b^{2}fr\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(-3\delta_{1}+4\delta_{1}fr^{2}\coth\left(fr^{2}+h\right)+\delta_{2}r^{2}\right)$
$Z_{7}=2b^{4}r\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)-8b^{4}fr^{3}\left(\delta_{2}r^{2}-2\delta_{1}\right)\coth\left(fr^{2}+h\right)\text{csch}^{4}\left(fr^{2}+h\right)$
$+2b^{2}\text{csch}^{2}\left(fr^{2}+h\right)\left(-8\delta_{1}f^{2}r^{3}\text{csch}^{2}\left(fr^{2}+h\right)+8\delta_{1}fr\coth\left(fr^{2}+h\right)+2\delta_{2}r\right)$
$Z_{8}=-8b^{2}f\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(-3\delta_{1}+4\delta_{1}fr^{2}\coth\left(fr^{2}+h\right)+\delta_{2}r^{2}\right)$
$Z_{9}=32b^{2}f^{2}r^{2}\coth^{2}\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(-3\delta_{1}+4\delta_{1}fr^{2}\coth\left(fr^{2}+h\right)+\delta_{2}r^{2}\right)$
$Z_{10}=10b^{4}\delta_{1}r^{2}\text{csch}^{4}\left(fr^{2}+h\right)+2b^{4}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)-32b^{4}\delta_{2}fr^{4}\coth\left(fr^{2}+h\right)\text{csch}^{4}\left(fr^{2}+h\right)$
$Z_{11}=16b^{4}f^{2}r^{4}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{6}\left(fr^{2}+h\right)+64b^{4}f^{2}r^{4}\left(\delta_{2}r^{2}-2\delta_{1}\right)\coth^{2}\left(fr^{2}+h\right)\text{csch}^{4}\left(fr^{2}+h\right)$
$\hskip 2.84544pt+40b^{4}fr^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\coth\left(fr^{2}+h\right)\text{csch}^{4}\left(fr^{2}+h\right)+16b^{2}f^{2}r^{2}\text{csch}^{4}\left(fr^{2}+h\right)\left(4b\delta_{1}r^{2}\coth\left(fr^{2}+h\right)-3\delta_{1}+\delta_{2}r^{2}\right)$
$Z_{12}=16b^{2}fr\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(-8\delta_{1}f^{2}r^{3}\text{csch}^{2}\left(br^{2}+c\right)+8\delta_{1}fr\coth\left(fr^{2}+h\right)+2\delta_{1}r\right)$
$Z_{13}=2b^{2}\text{csch}^{2}\left(fr^{2}+h\right)\left(2\delta_{2}-40\delta_{2}f^{2}r^{2}\text{csch}^{2}\left(fr^{2}+h\right)+8\delta_{1}f\coth\left(fr^{2}+f\right)\right)$
$Z_{14}=b^{4}r^{2}\left(\delta_{2}r^{2}-2\delta_{1}\right)\text{csch}^{4}\left(fr^{2}+h\right)+\delta_{2}+\phi_{1}$
$Z_{15}=8b^{2}f^{2}r^{4}\text{csch}^{4}\left(fr^{2}+h\right)+16b^{2}f^{2}r^{4}\coth^{2}\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)+2b^{2}\text{csch}^{2}\left(fr^{2}+h\right)-20b^{2}fr^{2}\coth\left(fr^{2}+h\right)$
×$\hskip 2.84544pt\times\text{csch}^{2}\left(fr^{2}+h\right)$
$F_{g}=-\frac{bB_{1}Z_{1}\phi_{22}\left(e^{fr^{2}+h}+1\right)}{2f\left(e^{fr^{2}+h}-1\right)Z_{4}}$
$F_{h}=\frac{aB_{1}\delta_{1}\phi_{22}\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{2}\phi_{5}}$
$+\frac{\delta_{1}\left(2b^{2}r\text{csch}^{2}\left(fr^{2}+h\right)-4b^{2}br^{3}\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\right)\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)}{r^{2}\phi_{7}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}$
$-\frac{\delta_{1}\left(4frZ_{4}\log(e)e^{2\left(fr^{2}+h\right)}+Z_{3}+Z_{5}\right)}{r^{2}\phi_{3}\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)}-\frac{2\delta_{1}}{r^{3}}+Z_{2}+\phi_{18}+\phi_{24}$
$F_{a}=\frac{2}{r}\left(\frac{2a^{2}W_{2}\text{csch}^{2}\left(br^{2}+c\right)}{2\phi_{7}\left(e^{br^{2}+c}-1\right)^{2}\left(e^{br^{2}+c}+1\right)^{2}}+A-\frac{r\phi_{24}}{2}+W_{1}\right)$
$\rho{\prime\prime}=\frac{-Z_{15}\left(a^{4}r^{2}\left(\text{$\delta$2}r^{2}-2\text{$\delta$1}\right)\text{csch}^{4}\left(br^{2}+c\right)+\text{$\delta$2}+\phi_{1}\right)}{\phi_{15}^{7}}$
$-\frac{3Z_{14}\left(2a^{2}r\text{csch}^{2}\left(br^{2}+c\right)-4a^{2}br^{3}\coth\left(br^{2}+c\right)\text{csch}^{2}\left(br^{2}+c\right)\right)^{2}}{\phi_{15}^{4}}$
$-\frac{2\left(Z_{6}+Z_{7}\right)\left(2a^{2}r\text{csch}^{2}\left(br^{2}+c\right)-4a^{2}br^{3}\coth\left(br^{2}+c\right)\text{csch}^{2}\left(br^{2}+c\right)\right)}{\phi_{15}^{7}}$
$+\frac{Z_{8}+Z_{9}+Z_{10}-Z_{11}-Z_{12}+Z_{13}+Z_{14}}{2\phi_{15}^{6}}$
$T_{1}=b^{3}\left(e^{2\left(fr^{2}+h\right)}-1\right)^{2}\text{csch}^{4}\left(fr^{2}+h\right)\left(2A_{1}f+bB_{1}\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)\right)^{2}$
$T_{2}=2A_{1}f\left(e^{2\left(fr^{2}+h\right)}+1\right)+2bB_{1}e^{fr^{2}+h}+bB_{1}\left(e^{2\left(fr^{2}+h\right)}+1\right)\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)$
$T_{3}=2bf\phi_{2}\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(8bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$T_{4}=fr^{2}\log(e)\left(2A_{1}f\left(e^{2\left(fr^{2}+h\right)}+1\right)+2bB_{1}e^{fr^{2}+h}+bB_{1}\left(e^{2\left(fr^{2}+h\right)}+1\right)\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)\right)+\phi_{2}$
$T_{5}=\text{csch}^{2}\left(fr^{2}+h\right)\left(b^{2}r^{2}\left(b^{2}-4f^{2}r^{2}\right)\text{csch}^{2}\left(fr^{2}+h\right)+5b^{2}-4f^{2}r^{2}\right)$
$+8f^{2}r^{2}\coth^{2}\left(fr^{2}+h\right)\left(b^{2}r^{2}\text{csch}^{2}\left(fr^{2}+h\right)-1\right)+2f\coth\left(fr^{2}+h\right)\left(5-3b^{2}r^{2}\text{csch}^{2}\left(fr^{2}+h\right)\right)$
$D_{1}=b^{4}\delta_{2}r^{4}\phi_{5}\text{csch}^{4}\left(fr^{2}+h\right)-16b^{3}B_{1}\delta_{1}f^{2}r^{4}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}\text{csch}^{2}\left(fr^{2}+h\right)$
$+4b^{2}\delta_{1}fr^{2}\left(e^{2\left(fr^{2}+h\right)}-1\right)\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(4bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$+8bb^{2}B_{1}\delta_{1}fr^{2}\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}\text{csch}^{2}\left(fr^{2}+h\right)-2b^{2}\phi_{5}\left(\delta_{1}-\delta_{2}r^{2}\right)\text{csch}^{2}\left(fr^{2}+h\right)$
$-16bB_{1}\delta_{1}f^{2}r^{2}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}+16bB_{1}\delta_{1}f\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}+\delta_{2}\phi_{5}$
$D_{2}=4b^{5}B_{1}\delta_{2}fr^{5}\log(e)\left(-e^{fr^{2}+h}\right)\left(e^{fr^{2}+h}+1\right)e^{fr^{2}+h}\text{csch}^{4}\left(fr^{2}+h\right)+8b^{4}\delta_{2}fr^{5}\phi_{2}\log(e)e^{2\left(fr^{2}+h\right)}\text{csch}^{4}\left(fr^{2}+h\right)$
$-8b^{4}\delta_{2}fr^{5}\phi_{5}\coth\left(fr^{2}+h\right)\text{csch}^{4}\left(fr^{2}+h\right)+4b^{4}\delta_{2}r^{3}\phi_{5}\text{csch}^{4}\left(fr^{2}+h\right)$
$-64b^{3}B_{1}\delta_{1}f^{3}rr^{4}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)$
$-32b^{3}B_{1}\delta_{1}f^{2}rr^{2}\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)$
$+16b^{2}\delta_{1}f^{2}r^{3}\left(e^{2\left(fr^{2}+h\right)}-1\right)\coth^{2}\left(fr^{2}+h\right)\text{csch}^{3}\left(fr^{2}+h\right)\left(4bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$-8b^{2}fr\phi_{5}\left(\delta_{1}-\delta_{2}r^{2}\right)\coth\left(fr^{2}+h\right)\text{csch}^{2}\left(fr^{2}+h\right)-64bB_{1}\delta_{1}f^{3}r^{3}\log^{3}(e)e^{3\left(br^{2}+c\right)}$
$-32bB_{1}\delta_{1}f^{3}r^{3}\log^{3}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}-32bB_{1}\delta_{1}f^{2}r\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}$
$+32bB_{1}\delta_{1}f^{2}r\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}+64bB_{1}\delta_{1}f^{2}r\log^{2}(e)e^{3\left(fr^{2}+h\right)}$
$+4bB_{1}\delta_{2}fr\log(e)\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)e^{fr^{2}+h}+8b\delta_{2}r\phi_{2}\log(e)e^{2\left(fr^{2}+h\right)}$
$D_{3}=8\delta_{1}f^{2}r^{3}\log(e)e^{fr^{2}+h}\left(e^{2\left(fr^{2}+h\right)}-1\right)\coth\left(fr^{2}+h\right)$
$D_{4}=D_{3}\left(2A_{1}fe^{fr^{2}+h}+bB_{1}e^{fr^{2}+h}\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)+2bB_{1}fr^{2}\log(e)+3B_{1}b\right)$
$-16bB_{1}\delta_{1}f^{3}r^{5}\log^{3}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}-32bB_{1}\delta_{1}f^{3}r^{5}\log^{3}(e)e^{3\left(fr^{2}+h\right)}$
$+8\delta_{1}f^{2}r^{3}\log(e)e^{2\left(fr^{2}+h\right)}\coth\left(fr^{2}+h\right)\left(4bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$-4\delta_{1}f^{2}r^{3}\left(e^{2\left(fr^{2}+h\right)}-1\right)\text{csch}^{2}\left(fr^{2}+h\right)\left(4bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$-32bB_{1}\delta_{1}f^{2}r^{3}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}+1\right)e^{fr^{2}+h}+8bB_{1}\delta_{1}f^{2}r^{3}\log^{2}(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}$
$+16bB_{1}\delta_{1}f^{2}r^{3}\log^{2}(e)e^{3\left(fr^{2}+h\right)}+4\delta_{1}fr\left(e^{2\left(fr^{2}+h\right)}-1\right)\coth\left(fr^{2}+h\right)\left(4bB_{1}fr^{2}\log(e)e^{fr^{2}+h}+\phi_{2}\right)$
$+8bB_{1}\delta_{1}fr\log(e)\left(e^{2\left(fr^{2}+h\right)}-1\right)e^{fr^{2}+h}+4bB_{1}fr\log(e)\left(\delta_{2}r^{2}-\delta_{1}\right)\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)e^{fr^{2}+h}$
$+8fr\phi_{2}\log(e)\left(\text{$\delta$2}r^{2}-\text{$\delta$1}\right)e^{2\left(br^{2}+c\right)}+2\delta_{2}r\phi_{5}$
$D_{5}=-4b^{4}r^{2}-2\left(b^{2}\left(8f^{2}r^{4}+5\right)-4f^{2}r^{2}\right)\cosh\left(2\left(fr^{2}+h\right)\right)+12b^{2}fr^{2}\sinh\left(2\left(fr^{2}+h\right)\right)+10b^{2}$
$-5f\sinh\left(4\left(br^{2}+c\right)\right)+4f^{2}r^{2}\cosh\left(4\left(fr^{2}+h\right)\right)-12f^{2}r^{2}+10f\sinh\left(2\left(fr^{2}+h\right)\right)$
$D_{6}=4b^{2}r\left(e^{fr^{2}+h}-1\right)\left(e^{fr^{2}+h}+1\right)\left(1-2fr^{2}\coth\left(fr^{2}+h\right)\right)\text{csch}^{2}\left(fr^{2}+h\right)$
$V^{2}_{r}=-\frac{\phi_{15}\sinh^{2}\left(fr^{2}+h\right)\left(-8b^{2}fB_{1}T_{4}\log(e)e^{fr^{2}+h}\text{csch}^{2}\left(fr^{2}+h\right)-8b^{2}B_{1}T_{2}\log^{2}(e)e^{fr^{2}+h}+T_{1}+T_{3}\right)}{aT_{5}\left(e^{fr^{2}+h}-1\right)^{2}\left(e^{fr^{2}+h}+1\right)^{2}\left(fB_{1}\log\left(\frac{e^{fr^{2}+h}-1}{e^{fr^{2}+h}+1}\right)+2A_{1}b\right)^{2}}$
$V^{2}_{t}=\frac{\phi_{3}\left(e^{\text{bf}r^{2}+h}-1\right)\left(e^{\text{bf}r^{2}+h}+1\right)\sinh^{6}\left(\text{bf}r^{2}+h\right)\left(2b^{2}D_{3}\text{csch}^{2}\left(\text{bf}r^{2}+h\right)+D_{2}\right)}{b^{2}\delta_{1}D_{5}r\left(e^{2\left(\text{bf}r^{2}+h\right)}-1\right)^{3}\left(2A_{1}f+bB_{1}\log\left(\frac{e^{\text{bf}r^{2}+h}-1}{e^{\text{bf}r^{2}+h}+1}\right)\right){}^{2}}$
$-4D_{1}fr\phi_{15}\log(e)\left(e^{\text{bf}r^{2}+h}+1\right)e^{\text{bf}r^{2}+h}\sinh^{6}\left(\text{bf}r^{2}+h\right)\left(2A_{1}f+bB_{1}\log\left(\frac{e^{\text{bf}r^{2}+h}-1}{e^{\text{bf}r^{2}+h}+1}\right)\right)$
$-D_{6}D_{1}\sinh^{6}\left(\text{bf}r^{2}+h\right)\left(2A_{1}f+bB_{1}\log\left(\frac{e^{\text{bf}r^{2}+h}-1}{e^{\text{bf}r^{2}+h}+1}\right)\right)-4bB_{1}D_{1}fr\phi_{15}\log(e)e^{\text{bf}r^{2}+h}\sinh^{6}\left(\text{bf}r^{2}+h\right)$
$-4D_{1}fr\phi_{3}\log(e)\left(e^{\text{bf}r^{2}+h}-1\right)e^{\text{bf}r^{2}+h}\sinh^{6}\left(\text{bf}r^{2}+h\right)$